The Miller Modigliani theorem posits that debt policy is irrelevant, when it comes to firm value. Assume that you have a firm that is funded entirely with equity and has a beta (unlevered) of 0.90, the risk-free rate is 3% and the equity risk premium is 6%. What will happen to the cost of capital, if the firm moves to a 30% debt ratio?

a.) none is true

b.) the cost of capital will go up

c.) the cost of capital will remain unchnaged

d.) the cost of capital will go down
Answer #1

Beta Unlevered =0.90
New debt ratio =30
So debt equity ratio =3/7
when debt policy is irrelevant for firm value .the tax rate =0%
Beta levered =Beta Unlevered*(1+Debt/Equity) =0.90*(1+30/70) =1.2857

Higher the beta, higher is the cost of capital.

Option b is correct.the cost of capital will go up